# Panoramic Distortion-Aware Tokenization for Person Detection and Localization in Overhead Fisheye Images

### v3

@@ 1 Introduction
Figure 1: Our method converts an overhead fisheye image into an equirectangular image by remapping. To leverage the significance areas of smaller persons, a distortion-aware tokenization process based on a self-similar figure using a unit figure, which is colored dark red and consists of a square and two rectangles. Our transformer-based detector uses the resulting tokens to detect bounding boxes within the panoramic image, and then the boxes project to the fisheye image. The input image is taken from [Yang_ICCV_2023].
@@
Seidel et al[Seidel_VISAPP_2019] VISAPP’19 YOLOv2 [Redmon_CVPR_2017]     Orthogonal rectangle
Li et al[Li_AVSS_2019] AVSS’19 YOLOv3 [Redmon_arXiv_2018] Rotated rectangle
Tamura et al[Tamura_WACV_2019] WACV’19 YOLOv2 [Redmon_CVPR_2017] Rotated rectangle
RAPiD [Duan_CVPRW_2020] CVPRW’20 FPN [Lin_CVPR_2017] Rotated rectangle
OARPD [Qiao_MTA_2024] MTA’24 YOLOv8m [Varghese_ADICS_2024] Rotated rectangle
@@
Yang et al[Yang_ICCV_2023] ICCV’23    DAB-DETR [Liu_ICLR_2022] Rotated rectangle
Ours DAB-DETR [Liu_ICLR_2022] Rotated trapezoid       ✓       ✓
@@
Object detection methods are used in a wide range of applications in both the commercial and industrial fields, including visual surveillance [Fu_CVPR_2021, Lin_TIP_2012, Xie_ICPR_2010, Yuan_CVPR_2024], pedestrian detection [Dollar_CVPR_2009, Khan_CVPR_2023, Zhang_PAMI_2018], and robotics [Lee_ICRA_2018, Yoshimi_IROS_2006, Zhang_IROS_2024]. In particular, the interest in fisheye image-based person detection has been growing because of its varied applications, which include monitoring of public transportation and infrastructure, worker movement optimization in factories, and player movement analysis for use in sports strategy. Overhead fisheye cameras are suitable for use in these applications because they are omnidirectional and have a larger field of view than conventional narrow-view cameras; i.e., one overhead fisheye camera can substitute for several narrow-view cameras. However, the overhead fisheye images do not match the inputs of the detectors [Ghiasi_CVPR_2021, Huang_2025_CVPR, Wang_CVPR_2023, Zhang_2025_CVPR, Zong_ICCV_2023], which require perspective images. This mismatch is derived from person rotation and from the presence of small persons in the images. The person rotation factor leads to individual persons having various shapes and appearances in the fisheye images. Within the near image circle of circumferential fisheye images, smaller people are captured at a long distance from the fisheye cameras. Although person detection methods based on use of narrow-view cameras are well-established, deriving accurate techniques for person detection and localization from overhead fisheye images has remained an open challenge.
Person detection methods have been proposed that used convolutional neural networks (CNNs) [Krizhevsky_NIPS_2012] and transformers [Vaswani_NIPS_2017] in overhead fisheye images. Early CNN-based methods [Li_AVSS_2019, Seidel_VISAPP_2019] used detectors that were trained on perspective images. To improve the performance of the CNN-based methods [Duan_CVPRW_2020, Qiao_MTA_2024, Tamura_WACV_2019] and a transformer-based method [Yang_ICCV_2023], their training was also conducted using overhead fisheye images. However, it is difficult for these methods to detect persons accurately in overhead fisheye images because of the rotation and small-person problems.
@@
To investigate the effectiveness of these proposed methods, we conducted extensive experiments on three large-scale datasets [Duan_CVPRW_2020, Tezcan_WACV_2022, Yang_ICCV_2023]. This evaluation demonstrated that our proposed method outperforms both the conventional CNN-based [Duan_CVPRW_2020, Li_AVSS_2019, Qiao_MTA_2024, Seidel_VISAPP_2019, Tamura_WACV_2019] and transformer-based [Yang_ICCV_2023] methods. The major contributions of our study are summarized as follows:
@@ 2 Related work
Overhead-fisheye-image person detection. Person detection techniques in overhead fisheye images have been developed using various datasets [Blanco_SPIC_2021, Duan_CVPRW_2020, Li_AVSS_2019, Tezcan_WACV_2022, Yang_ICCV_2023]. Existing person detection methods for overhead fisheye images can be classified into two categories: CNN-based and transformer-based methods, as listed in Table 1. Seidel et al[Seidel_VISAPP_2019] proposed a pioneering CNN-based detection method. The method divides a fisheye image into upright image patches that are undistorted using the given camera parameters. Although these patches are perspective images, the method requires camera calibration before detection. Li et al[Li_AVSS_2019] proposed a CNN-based method to improve detection accuracy using focal windows that can extract upright images. Tamura et al[Tamura_WACV_2019] proposed a method that uses YOLOv2 [Redmon_CVPR_2017], which had been trained on perspective images from the COCO dataset [Lin_ECCV_2014] using rotation augmentation. To train a detector called RAPiD, Duan et al[Duan_CVPRW_2020] introduced the angle-aware loss function. Inspired by RAPiD being trained on fisheye images, OARPD [Qiao_MTA_2024] was proposed with use of the center distance intersection over union approach to address rotated bounding boxes. Yang et al[Yang_ICCV_2023] proposed a pioneering transformer-based method with rotation-equivariant training. For a query-based detector, this training strategy involves use of an additional query of a $90^{\circ}$-rotated image to tackle rotated persons. Both the CNN-based and transformer-based methods described above focused primarily on person rotation. Although the CNN-based and transformer-based methods can detect persons, images containing persons of various sizes cause performance degradation.
@@
Small object detection. Object detection accuracy tends to decrease for smaller objects. Small objects are often captured at long distances, e.g., in a dataset acquired using drones [Zhu_arXiv_2018]. Small object detection methods are generally classified into four categories: division-based methods, augmentation-based methods, enlargement-based methods, and spectrum-based methods. The division-based methods involve multi-scale training performed using image pyramids [Singh_CVPR_2018], image cropping [Singh_NIPS_2018], and feature pyramids [Gong_WACV_2021]. The augmentation-based methods [Ghiasi_CVPR_2021, Yu_WACV_2020] were proposed to realize data invariance. The enlargement-based methods enlarge the input images [Liu_TIP_2025, Unel_CVPRW_2019, Wang_PAMI_2021] and conduct super-resolution processing [Ji_ICPR_2021, Shermeyer_CVPRW_2019]. The spectrum-based methods enhance the discrimination of objects from a frequency perspective [Sun_2025_CVPR]. The methods noted above do not use the prior knowledge that smaller persons are only located near the image circles in overhead fisheye images. Therefore, these conventional methods do not work effectively for person detection in overhead fisheye images.
@@
Transformer tokenization. Dosovitskiy et al[Dosovitskiy_ICLR_2021] proposed a pioneering transformer-based method using a tokenization approach that divides an input image into square patches. Numerous tokenization methods have been proposed to date because these square patches lead to high computational costs. Window-based methods focus on the token groups using shifted windows [Liu_ICCV_2021], cross-shaped windows [Dong_CVPR_2022], dilated windows [Tu_ECCV_2022, Wang_PAMI_2024], and scalable sliding windows [Hassani_CVPR_2023]. In contrast to static window shapes, deformable attention methods [Xia_CVPR_2022] were proposed to improve flexibility. To address the problem of multiple tokens being generated in unimportant areas, e.g., in the background, DynamicViT [Rao_NIPS_2021] and BiFormer [Zhu_CVPR_2023] prune the areas, and SG-Former [Ren_ICCV_2023] aggregates the areas. Although the methods described above are all well-developed, these methods were designed for use with perspective images. For fisheye image applications, DarSwin [Athwale_ICCV_2023], which uses radial patches formed by polar partitioning, was proposed. However, this partitioning approach is not suitable for overhead fisheye images because a small number of the relevant tokens are generated in near image circles.
@@
Visual localization. Various localization methods have been established to determine either camera poses or person locations from an image. These methods can be classified into three categories: query-based methods, regression-based methods, and geometry-based methods. The query-based methods estimate camera poses using databases composed of geo-tagged images [Aiger_ICCV_2023, Balntas_ECCV_2018, Hyeon_ICCV_2021, Sarlin_CVPR_2021, Torii_PAMI_2021, YangL_ICCV_2019] and 3D scene structure models [Germain_CVPR_2021, Sarlin_CVPR_2019]. Before query-based methods can be used, the databases and models must first be built in a time-consuming process. The regression-based methods [Kendall_ICCV_2015, Walch_ICCV_2017] can estimate camera poses without the need for databases and models; however, their inferences can only be applied to the same scenes that were used for training. The geometry-based methods estimate horizontal distances of person positions based on the bounding boxes [Yang_ICCV_2023, Zhu_JVCIR_2019] and head points [DelBlanco_SPIC_2021] in overhead fisheye images. These geometry-based methods estimate person locations from image coordinates using camera parameters.
@@ 3.1 Geometric analysis of overhead fisheye images
We analyzed the LOAF training set [Yang_ICCV_2023] to understand the unique characteristics of overhead fisheye images, which differ from perspective images. We remapped overhead fisheye images to equirectangular images, as shown in Figure 2(a). A downward-looking fisheye camera observes the lower hemisphere, with incident angles $\theta$ ranging from 0 (nadir) to 90 (horizon). Figure 2(b) shows the distribution of bounding box heights and widths along different incident angles $\theta$. Our analysis revealed two key findings:
@@
Note that the bias in the second finding has two main causes. First, peripheral portions at large incident angles subtend larger ground areas. Second, due to motion parallax [Gibson_JEP_1959, Hartley_CUP_2004], people farther from the camera exhibit smaller apparent image motion and remain in view longer, while nearer people traverse the field of view more quickly.
@@ 3.2 Design of our proposed method
We designed our proposed method based on the two key findings, as described in Section 3.1, because conventional detection methods, which assume uniform object distribution, may not be optimal for overhead fisheye images. The first finding suggests that our method uses the prior knowledge of the linear change in predictable bounding box sizes. The second finding indicates that focusing on bounding boxes with large incident angles is effective in terms of the number of bounding boxes. Conventional detectors allocate computation approximately uniformly across the image and do not explicitly exploit $\theta$-dependent scale or distribution biases. Therefore, we designed our tokenization method to specifically account for these geometric relationships, focusing on the portions with large incident angles.
@@ 3.3 Panoramic distortion-aware tokenization
Each of these square tiles contains $a_{k}\times a_{k}$ feature vectors, where $a_{k}$ represents the number of feature vectors along each side of a tile in the $k$th region, as shown in Figure 3(b). Our tiling is defined on the basis of the self-similar figures when using $K$ regions, with the exception that the top two rows in region 1 have tiles of the same size due to finite divisions. The unit figure consists of a square and two rectangles, and each rectangle can be divided recursively using another set composed of a square and two rectangles, as illustrated in the rightmost part of Figure 3(b). Note that our division method was inspired by the Cantor ternary set [Smith_LMS_1874], but our self-similar figure is not a fractal according to Mandelbrot’s definition [Mandelbrot_Science_1967].
Although the tile sizes vary among regions 3, 4, and 5, as shown in Figure 3(a), the person-bounding boxes appear similar in size relative to their respective tiles. This design maintains similar complexity across all regions, despite the varying absolute sizes of persons in the panoramic image. As illustrated in Figure 3(b), our tokenization method uses recursive divisions to maintain consistent person-to-tile size ratios, particularly in portions with large incident angles.
@@ 3.4 Network architecture
Our network architecture consists of four main components: panoramic image conversion, a detector head based on the dynamic anchor box detection with transformers (DAB-DETR) [Liu_ICLR_2022], a backbone based on SG-Former [Ren_ICCV_2023], and our proposed PDAT.
Panoramic image conversion. Our method converts an input fisheye image into an equirectangular panoramic image via remapping, as illustrated in Figure 4. When the input image is rectangular, we center-crop the image. To address any uncalibrated cameras, we assume that the input images represent the stereographic projection, which is a standard fisheye camera model. In the case of radial distortion [Wakai_ECCV_2022], this projection uses a trigonometric function: $\gamma=2f_{w}\cdot\tan(\theta/2)$, where $\gamma$ is the distortion, $f_{w}$ is the focal length, and $\theta$ represents the incident angle, as shown in Figure 2(a). We can then determine the focal length by fitting the image circle using $90^{\circ}$ incident angles. The image coordinates in the fisheye images are projected onto a unit sphere in world coordinates using backprojection [Wakai_ECCV_2022], and then the world coordinates are projected onto an equirectangular image with an arbitrary image width [Sharma_AIVR_2019]. We use panoramic images with incident angles from $0^{\circ}$ to $90^{\circ}$ because the incident angle range from $90^{\circ}$ to $180^{\circ}$ represents the directions behind the cameras.
Detector and backbone selection. Following the approach in [Yang_ICCV_2023], we use DAB-DETR [Liu_ICLR_2022], which has shown strong object detection performances. Window-based transformers are not suitable for use in the backbone of DAB-DETR because PDAT divides the feature maps nonuniformly. Therefore, we use an aggregation-based transformer called SG-Former [Ren_ICCV_2023] for the backbone. Note that SG-Former is the only aggregation-based method of this type available. The proposed network architecture is shown in Figure 4. The detector head block represents DAB-DETR without the backbone, and the remaining blocks represent SG-Former and PDAT. The patch embedding is extracted using the CNN, and the transformer block is a vanilla transformer block [Dosovitskiy_ICLR_2021]. The implementations of DAB-DETR and SG-Former follow the details provided in their respective papers. We explain the integration of PDAT into SG-Former in the following two steps: 1) SG-Former attention mechanism and 2) PDAT integration details.
@@ 3.5 Training and inference
Training. We generate equirectangular images from each training set. In this generation procedure, we avoid splitting the bounding boxes in the vertical image edges by changing the definition of the 0 azimuth angles. We then train our network using the panoramic images by following [Liu_ICLR_2022].
Inference. First, we convert an input fisheye image into a panoramic image by a remapping process, as illustrated in Figure 4. Second, the remapped panoramic image is then fed into our network. Finally, we project the detected bounding boxes onto the fisheye image, on which the shape of each bounding box is a rotated trapezoid. This projection represents the backprojection from a fisheye image to a panoramic image.

### v4

@@ 1 Introduction
Figure 1: Our method converts an overhead fisheye image into an equirectangular image by remapping. To leverage the significance areas of smaller persons, a distortion-aware tokenization process based on a self-similar figure using a unit figure, which is colored dark red and consists of a square and two rectangles. Our transformer-based detector uses the resulting tokens to detect bounding boxes within the panoramic image, and then the boxes project to the fisheye image. The input image is taken from [73].
@@
Seidel et al. [48] VISAPP’19 YOLOv2 [41]     Orthogonal rectangle
Li et al. [26] AVSS’19 YOLOv3 [42] Rotated rectangle
Tamura et al. [55] WACV’19 YOLOv2 [41] Rotated rectangle
RAPiD [11] CVPRW’20 FPN [28] Rotated rectangle
OARPD [39] MTA’24 YOLOv8m [60] Rotated rectangle
@@
Yang et al. [73] ICCV’23    DAB-DETR [31] Rotated rectangle
Ours DAB-DETR [31] Rotated trapezoid       ✓       ✓
@@
Object detection methods are used in a wide range of applications in both the commercial and industrial fields, including visual surveillance [12, 27, 71, 76], pedestrian detection [8, 23, 79], and robotics [25, 74, 78]. In particular, the interest in fisheye image-based person detection has been growing because of its varied applications, which include monitoring of public transportation and infrastructure, worker movement optimization in factories, and player movement analysis for use in sports strategy. Overhead fisheye cameras are suitable for use in these applications because they are omnidirectional and have a larger field of view than conventional narrow-view cameras; i.e., one overhead fisheye camera can substitute for several narrow-view cameras. However, the overhead fisheye images do not match the inputs of the detectors [14, 19, 69, 77, 83], which require perspective images. This mismatch is derived from person rotation and from the presence of small persons in the images. The person rotation factor leads to individual persons having various shapes and appearances in the fisheye images. Within the near image circle of circumferential fisheye images, smaller people are captured at a long distance from the fisheye cameras. Although person detection methods based on use of narrow-view cameras are well-established, deriving accurate techniques for person detection and localization from overhead fisheye images has remained an open challenge.
Person detection methods have been proposed that used convolutional neural networks (CNNs) [24] and transformers [61] in overhead fisheye images. Early CNN-based methods [26, 48] used detectors that were trained on perspective images. To improve the performance of the CNN-based methods [11, 39, 55] and a transformer-based method [73], their training was also conducted using overhead fisheye images. However, it is difficult for these methods to detect persons accurately in overhead fisheye images because of the rotation and small-person problems.
@@
To investigate the effectiveness of these proposed methods, we conducted extensive experiments on three large-scale datasets [11, 56, 73]. This evaluation demonstrated that our proposed method outperforms both the conventional CNN-based [11, 26, 39, 48, 55] and transformer-based [73] methods. The major contributions of our study are summarized as follows:
@@ 2 Related work
Overhead-fisheye-image person detection. Person detection techniques in overhead fisheye images have been developed using various datasets [6, 11, 26, 56, 73]. Existing person detection methods for overhead fisheye images can be classified into two categories: CNN-based and transformer-based methods, as listed in Table 1. Seidel et al. [48] proposed a pioneering CNN-based detection method. The method divides a fisheye image into upright image patches that are undistorted using the given camera parameters. Although these patches are perspective images, the method requires camera calibration before detection. Li et al. [26] proposed a CNN-based method to improve detection accuracy using focal windows that can extract upright images. Tamura et al. [55] proposed a method that uses YOLOv2 [41], which had been trained on perspective images from the COCO dataset [29] using rotation augmentation. To train a detector called RAPiD, Duan et al. [11] introduced the angle-aware loss function. Inspired by RAPiD being trained on fisheye images, OARPD [39] was proposed with use of the center distance intersection over union approach to address rotated bounding boxes. Yang et al. [73] proposed a pioneering transformer-based method with rotation-equivariant training. For a query-based detector, this training strategy involves use of an additional query of a $90^{\circ}$-rotated image to tackle rotated persons. Both the CNN-based and transformer-based methods described above focused primarily on person rotation. Although the CNN-based and transformer-based methods can detect persons, images containing persons of various sizes cause performance degradation.
@@
Small object detection. Object detection accuracy tends to decrease for smaller objects. Small objects are often captured at long distances, e.g., in a dataset acquired using drones [82]. Small object detection methods are generally classified into four categories: division-based methods, augmentation-based methods, enlargement-based methods, and spectrum-based methods. The division-based methods involve multi-scale training performed using image pyramids [51], image cropping [52], and feature pyramids [16]. The augmentation-based methods [14, 75] were proposed to realize data invariance. The enlargement-based methods enlarge the input images [30, 59, 67] and conduct super-resolution processing [21, 50]. The spectrum-based methods enhance the discrimination of objects from a frequency perspective [54]. The methods noted above do not use the prior knowledge that smaller persons are only located near the image circles in overhead fisheye images. Therefore, these conventional methods do not work effectively for person detection in overhead fisheye images.
@@
Transformer tokenization. Dosovitskiy et al. [10] proposed a pioneering transformer-based method using a tokenization approach that divides an input image into square patches. Numerous tokenization methods have been proposed to date because these square patches lead to high computational costs. Window-based methods focus on the token groups using shifted windows [33], cross-shaped windows [9], dilated windows [58, 68], and scalable sliding windows [18]. In contrast to static window shapes, deformable attention methods [70] were proposed to improve flexibility. To address the problem of multiple tokens being generated in unimportant areas, e.g., in the background, DynamicViT [40] and BiFormer [81] prune the areas, and SG-Former [44] aggregates the areas. Although the methods described above are all well-developed, these methods were designed for use with perspective images. For fisheye image applications, DarSwin [2], which uses radial patches formed by polar partitioning, was proposed. However, this partitioning approach is not suitable for overhead fisheye images because a small number of the relevant tokens are generated in near image circles.
@@
Visual localization. Various localization methods have been established to determine either camera poses or person locations from an image. These methods can be classified into three categories: query-based methods, regression-based methods, and geometry-based methods. The query-based methods estimate camera poses using databases composed of geo-tagged images [1, 3, 20, 47, 57, 72] and 3D scene structure models [13, 46]. Before query-based methods can be used, the databases and models must first be built in a time-consuming process. The regression-based methods [22, 64] can estimate camera poses without the need for databases and models; however, their inferences can only be applied to the same scenes that were used for training. The geometry-based methods estimate horizontal distances of person positions based on the bounding boxes [73, 80] and head points [7] in overhead fisheye images. These geometry-based methods estimate person locations from image coordinates using camera parameters.
@@ 3.1 Geometric analysis of overhead fisheye images
We analyzed the LOAF training set [73] to understand the unique characteristics of overhead fisheye images, which differ from perspective images. We remapped overhead fisheye images to equirectangular images, as shown in Figure 2(a). A downward-looking fisheye camera observes the lower hemisphere, with incident angles $\theta$ ranging from 0 (nadir) to 90 (horizon). Figure 2(b) shows the distribution of bounding box heights and widths along different incident angles $\theta$. Our analysis revealed two key findings:
@@
Note that the bias in the second finding has two main causes. First, peripheral portions at large incident angles subtend larger ground areas. Second, due to motion parallax [15, 17], people farther from the camera exhibit smaller apparent image motion and remain in view longer, while nearer people traverse the field of view more quickly.
@@ 3.2 Design of our proposed method
We designed our proposed method based on the two key findings, as described in Section 3.1, because conventional detection methods, which assume uniform object distribution, may not be optimal for overhead fisheye images. The first finding suggests that our method uses the prior knowledge of the linear change in predictable bounding box sizes. The second finding indicates that focusing on bounding boxes with large incident angles is effective in terms of the number of bounding boxes. Conventional detectors allocate computation approximately uniformly across the image and do not explicitly exploit $\theta$-dependent scale or distribution biases. Therefore, we designed our tokenization method to specifically account for these geometric relationships, focusing on the portions with large incident angles. The detailed mathematical derivation of these relationships is provided in the supplementary material.
@@ 3.3 Panoramic distortion-aware tokenization
Each of these square tiles contains $a_{k}\times a_{k}$ feature vectors, where $a_{k}$ represents the number of feature vectors along each side of a tile in the $k$th region, as shown in Figure 3(b). Our tiling is defined on the basis of the self-similar figures when using $K$ regions, with the exception that the top two rows in region 1 have tiles of the same size due to finite divisions. The unit figure consists of a square and two rectangles, and each rectangle can be divided recursively using another set composed of a square and two rectangles, as illustrated in the rightmost part of Figure 3(b). Note that our division method was inspired by the Cantor ternary set [53], but our self-similar figure is not a fractal according to Mandelbrot’s definition [36].
Although the tile sizes vary among regions 3, 4, and 5, as shown in Figure 3(a), the person-bounding boxes appear similar in size relative to their respective tiles. This design maintains similar complexity across all regions, despite the varying absolute sizes of persons in the panoramic image. As illustrated in Figure 3(b), our tokenization method uses recursive divisions to maintain consistent person-to-tile size ratios, particularly in portions with large incident angles. The mathematical formulation of PDAT is given in the supplementary material.
@@ 3.4 Network architecture
Our network architecture consists of four main components: panoramic image conversion, a detector head based on the dynamic anchor box detection with transformers (DAB-DETR) [31], a backbone based on SG-Former [44], and our proposed PDAT.
Panoramic image conversion. Our method converts an input fisheye image into an equirectangular panoramic image via remapping, as illustrated in Figure 4. When the input image is rectangular, we center-crop the image. To address any uncalibrated cameras, we assume that the input images represent the stereographic projection, which is a standard fisheye camera model. In the case of radial distortion [62], this projection uses a trigonometric function: $\gamma=2f_{w}\cdot\tan(\theta/2)$, where $\gamma$ is the distortion, $f_{w}$ is the focal length, and $\theta$ represents the incident angle, as shown in Figure 2(a). We can then determine the focal length by fitting the image circle using $90^{\circ}$ incident angles. The image coordinates in the fisheye images are projected onto a unit sphere in world coordinates using backprojection [62], and then the world coordinates are projected onto an equirectangular image with an arbitrary image width [49]. We use panoramic images with incident angles from $0^{\circ}$ to $90^{\circ}$ because the incident angle range from $90^{\circ}$ to $180^{\circ}$ represents the directions behind the cameras.
Detector and backbone selection. Following the approach in [73], we use DAB-DETR [31], which has shown strong object detection performances. Window-based transformers are not suitable for use in the backbone of DAB-DETR because PDAT divides the feature maps nonuniformly. Therefore, we use an aggregation-based transformer called SG-Former [44] for the backbone. Note that SG-Former is the only aggregation-based method of this type available. The proposed network architecture is shown in Figure 4. The detector head block represents DAB-DETR without the backbone, and the remaining blocks represent SG-Former and PDAT. The patch embedding is extracted using the CNN, and the transformer block is a vanilla transformer block [10]. The implementations of DAB-DETR and SG-Former follow the details provided in their respective papers. We explain the integration of PDAT into SG-Former in the following two steps: 1) SG-Former attention mechanism and 2) PDAT integration details.
@@ 3.5 Training and inference
Training. We generate equirectangular images from each training set. In this generation procedure, we avoid splitting the bounding boxes in the vertical image edges by changing the definition of the 0 azimuth angles. We then train our network using the panoramic images by following [31].
Inference. First, we convert an input fisheye image into a panoramic image by a remapping process, as illustrated in Figure 4. Second, the remapped panoramic image is then fed into our network. Finally, we project the detected bounding boxes onto the fisheye image, on which the shape of each bounding box is a rotated trapezoid. This projection represents the backprojection from a fisheye image to a panoramic image. The supplementary material describes details of the training and inference procedures.